OpenStudy (anonymous):

Find all solutions to 2sin(x)cos(x) - sin(x)=0 in the interval [0,2pi), keep answer in radians with pi. Can you also explain to me how you got the answer please

OpenStudy (amoodarya):

2sin(x)cos(x) - sin(x)=0 sin(2x)-sin(x) =0 sin (2x)=sin (x) can you go on ?

OpenStudy (amoodarya):



jimthompson5910 (jim_thompson5910):

2sin(x)cos(x) - sin(x)=0 sin(x)(2cos(x) - 1) = 0 sin(x) = 0 or 2cos(x) - 1 = 0 sin(x) = 0 or 2cos(x) = 1 sin(x) = 0 or cos(x) = 1/2 I'll let you finish

OpenStudy (amoodarya):